OpenStudy (anonymous):

Vin=45V, R1=1.5Ω, R2=2Ω, R3=6Ω, A=2 and B=0.5 what is the RTH ?

OpenStudy (benzo):

sara do you mean I1= vin/(r3-b*r1*r3 vo=vin*(r3-(b*r1*r3)+r1+r2+a*r2)/(r3-b*r1*r3)

OpenStudy (anonymous):

Tell me the othersQ6 answews, pls.. I have the same values

OpenStudy (anonymous):

5 7.5 7.5

OpenStudy (anonymous):

She/He wrote that and its correct only if you have those values.

OpenStudy (anonymous):

I have the same values. and did you get the RTH?

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

@benzo 1-Vo* (R3-(B*R1*R3)+R1+R2+A*R2)/(R3-B*R1*R3)=VIN 2-i1=V0/(R3-B*R1*R3) 3- vo = vth this is the formula of q6 did you mean that ?

OpenStudy (anonymous):

do you guys have any news abt the rth?

OpenStudy (anonymous):

yeah....I saw sum1 had postd d formula... Its Vth/Isc... Bt don knw how to calculate Isc.. :P Can sumbdy help me out pls.?

OpenStudy (anonymous):

Rth=5 for ur values @EngSara

OpenStudy (anonymous):

Thx same for me

OpenStudy (anonymous):

thanks @1431mahi